# Discontinuous Transition to Active Nematic Turbulence

### v1

@@ Abstract
Active fluids exhibit chaotic flows at low Reynolds number known as active turbulence. Whereas the statistical properties of the chaotic flows are increasingly well understood, the nature of the transition from laminar to turbulent flows as activity increases remains unclear. Here, through simulations of a minimal model of unbounded active nematics, we find that the transition to active turbulence is discontinuous. We show that the transition features a jump in the mean-squared velocity, as well as bistability and hysteresis between laminar and chaotic flows. From distributions of finite-time Lyapunov exponents, we identify the transition at a value $A^{*}\approx 4900$ of the dimensionless activity number. Below the transition to chaos, we find subcritical bifurcations that feature bistability of different laminar patterns. These bifurcations give rise to oscillations and to chaotic transients, which become very long close to the transition to turbulence. Overall, our findings contrast with the continuous transition to turbulence in channel confinement, where turbulent puffs emerge within a laminar background. We propose that, without confinement, the long-range hydrodynamic interactions of Stokes flow suppress the spatial coexistence of different flow states, and thus render the transition discontinuous.
I Introduction
@@
How do laminar flows become turbulent? Scientists have been seeking answers to this deceptively simple question since the seminal experiments by Reynolds in 1883 1. Reynolds’ original work showed that the onset of turbulence is controlled by a dimensionless parameter — the Reynolds number — which compares inertial to viscous forces. Despite having a single control parameter, the nature of the transition to turbulence has been debated for over a century 2. Detailed experiments in the past decade showed that, in pipe and Couette flow, the onset of turbulence follows a continuous second-order phase transition, with critical exponents in the directed percolation universality class 3; 4; 5; 6; 7; 8; 9; 10; 11; 12. Even more recent work showed that, under body forces 13, the transition becomes discontinuous, similar to first-order phase transitions.
@@
In stark contrast to high-Reynolds-number inertial turbulence, two decades ago, turbulent-like flows were discovered in bacterial suspensions at low Reynolds number 14. Such spontaneous chaotic flows, now broadly known as active turbulence, have since been observed in a variety of (predominantly biological) fluids 15. In addition to bacterial suspensions 14; 16; 17; 18; 19; 20; 21, active turbulence is found in suspensions of microtubules and molecular motors 22; 23; 24; 25; 26; 27; 28, as well as in epithelial monolayers 29; 30; 31 and self-propelled particles 32. These active systems consume stored energy to power local internal driving 33. When this internal driving is strong enough, the flows become chaotic. But how and when does chaos emerge?
@@
For active polar fluids, simulations of multiple models showed that chaos is reached through a sequence of oscillatory instabilities 34; 35; 36; 37; 15; 38. Other works have reported transitions to turbulence by varying either activity 39; 40; 41; 42 or swimmer concentration 43; 44. For active nematics, simulations showed that, in a channel, the system goes from simple shear to oscillatory flow and to a vortex chain before reaching turbulence45; 46; 47; 48; 49; 50. Similar states were seen experimentally in bacterial and in microtubule suspensions by increasing the confinement size 51; 25; 52. Importantly, simulations showed that the transition to chaos follows a scenario of directed percolation similar to pipe flow, with localised puffs of turbulence eventually spreading throughout the system as activity increases 47. In the absence of confinement, both simulations 53; 54; 55 and experiments 26 revealed a sequence of bend instabilities leading to turbulence. However, key questions remain: What is the nature of the bifurcations that end up in chaos? And what is the activity threshold, as well as the type, of the transition to turbulence?
@@
Here, we combine approaches from dynamical systems and statistical physics to investigate the transition to active nematic turbulence in a minimal model without confinement. We find that the transition is discontinuous, with the system exhibiting a jump in flow intensity and fluctuations, as well as a region of bistability and hysteresis, as in first-order phase transitions. We then build the bifurcation diagram of the initial instabilities that ultimately lead to the outbreak of chaos. Thereby, we expose the earliest appearances of subcritical bifurcations, oscillations, and chaotic transients that culminate in the discontinuous transition to turbulence. When compared to the directed-percolation scenario in channels 47, our findings suggest that confinement can change the nature of the transition from discontinuous to continuous. Therefore, our results limit the applicability of the notion of universality in the transition to active turbulence. Our results also imply that, when unconstrained, active nematic turbulence sets in at a value $A^{*}\approx 4900$ of the dimensionless activity number, which therefore plays a role similar to that of the critical Reynolds number for the transition to inertial turbulence.
II Active nematics hydrodynamics
@@
We study a minimal hydrodynamic model of incompressible active nematics in two dimensions 55. Neglecting inertia by taking zero Reynolds number, the force balance equation reads
$0=\partial_{\alpha}P+\partial_{\beta}\left(\sigma_{\alpha\beta}+\sigma^{\text{a}}_{\alpha\beta}\right).$ (1)
@@
Here, the symmetric part of the stress tensor is $\sigma_{\alpha\beta}=2\eta v_{\alpha\beta}-\zeta q_{\alpha\beta}$, with $\eta$ the shear viscosity, $v_{\alpha\beta}=1/2\left(\partial_{\alpha}v_{\beta}+\partial_{\beta}v_{\alpha}\right)$ the symmetric part of the strain rate tensor, $\zeta$ the active stress coefficient, and $q_{\alpha\beta}=n_{\alpha}n_{\beta}-1/2\delta_{\alpha\beta}$ the nematic orientation tensor defined by the director field $\bm{n}=(\cos\theta,\sin\theta)$. The director has unit norm as we consider the system to be deep in the nematic phase and consequently defect-free. Respectively, the antisymmetric stress is $\sigma^{\text{a}}_{\alpha\beta}=1/2\left(n_{\alpha}h_{\beta}-h_{\alpha}n_{\beta}\right)$, where the molecular field $h_{\alpha}=K\nabla^{2}n_{\alpha}$ is calculated from minimizing the Frank free energy for nematic elasticity in the one-constant approximation 56.
@@
where $\omega_{\alpha\beta}=1/2\left(\partial_{\alpha}v_{\beta}-\partial_{\beta}v_{\alpha}\right)$ is the vorticity tensor and $\gamma$ is the rotational viscosity. For simplicity, following Ref. 55, we ignore flow alignment; including it does not alter the scaling properties of the turbulent flows 57.
@@
$-\nabla^{4}\psi=\frac{1}{2}\frac{R}{A}\nabla^{4}\theta+S\left[\frac{1}{2}\left(\partial_{x}^{2}-\partial_{y}^{2}\right)\sin(2\theta)-\partial_{xy}^{2}\cos(2\theta)\right],$ (3)
where $S=\zeta/|\zeta|$ is the sign of the active stress taken to be $S=-1$, corresponding to contractile stress, $R=\gamma/\eta$ is the viscosity ratio taken to be $R=1$, and the activity number $A=L^{2}/\ell^{2}_{c}$ where $\ell_{c}=\sqrt{K/(|\zeta|R)}$ is the active length from balancing active and nematic elastic stresses. In the following sections, we vary $A$. Since the active time $\tau_{\text{a}}$ itself varies with activity, we use the nematic relaxation time $\tau_{\text{r}}=\gamma L^{2}/K$ as the time unit in the plots.
@@
Equations 3 and 4 provide a minimal hydrodynamic model that captures the scalings of the velocity power spectrum of fully-developed turbulent flows 55, which have been successfully compared to experiments 28.
III Transition to Active Turbulence
@@
Figure 1: Transition to active turbulence. 1-1, Snapshots of the flow field in the laminar (1), bistable (1,1), and chaotic (1) regimes. The background shows the nematic director through line integral convolution, color denotes the stream function $\psi$, and arrows depict the velocity. 1, Mean squared velocity (MSV) $\langle v^{2}\rangle_{\mathbf{r}}$ from multiple realizations at different activity number $A$. Gray points are individual realizations, and purple points are the ensemble averages. Regimes of laminar and chaotic flow are separated by a jump in MSV and a transition region where both states are possible. Labels 1-1 correspond to the snapshots above. 1, MSV from a single realization with increasing (red) or decreasing (blue) activity. The transition between laminar and turbulent flows exhibits hysteresis. Shaded regions show the s.d. of temporal fluctuations in MSV. 1, The maximal Lyapunov exponent (MLE) $\Lambda$ shows the transition from non-chaotic to chaotic dynamics as activity increases. 1, The fraction of time spent in the chaotic state $f_{\text{c}}$ also shows a transition from non-chaotic to chaotic dynamics with an intermediate region where both are possible. The inset zooms into the end of the transition region. By increasing the integration time $t_{\text{f}}$ (blue to red), we identify the transition to turbulence at $A^{*}\approx 4900$, above which the flow never laminarizes. 1, Exemplary time series of the stretching number $\alpha$ showing a chaotic transient ($\alpha\neq 0$) that eventually decays into a laminar state ($\alpha=0$). 1, Histogram of 1. 1-1, For increasing activity, the distributions of the stretching number show the change from completely laminar (1) to completely chaotic flow (1) through a transition region with a bimodal distribution indicative of bistability (1). The unit of time for all quantities in this figure is $\tau_{\text{r}}$.
@@
When and how do turbulent flows emerge? We first approach this question statistically. To find and characterize the transition to turbulence, we numerically integrate Eqs. 3 and 4 at different values of the activity number $A$, starting from a uniformly aligned director field (see methods in Section A.1). To obtain different realizations, we add noise for a short initial period before the evolution continues deterministically (see Section A.1 for details). Looking at many such realizations, we identify three broad regimes with qualitatively different dynamics: (i) At low activity, we find laminar flow states, such as vortex patterns, either steady or oscillating (Fig. 1 and Movie S1); (ii) at intermediate activity, both laminar flow patterns and chaos are possible (Figs. 1 and 1); and (iii) at high activity, we find chaotic flow (Fig. 1 and Movie S2).
III.1 Jump in flow intensity, bistability, and hysteresis
@@
To quantify the flow intensity in these different regimes, we compute the space-averaged mean squared velocity (MSV), $\langle v^{2}\rangle_{\mathbf{r}}$, for many realizations at each activity (Fig. 1). At low activity ($A\lesssim 3800$), the MSV increases linearly with activity. There is very little variation between realizations (grey points in Fig. 1; purple points show the ensemble average). At high activity, ($A\gtrsim 5000$), the MSV once again increases linearly, consistent with previous findings 58, though now with a steeper slope. In addition, the variation between realizations is much higher than in the laminar states at low activity. In between, we find a transition region ($3800\lesssim A\lesssim 5000$) showing bistability between laminar and chaotic flows: Both types of flow can occur at a given activity.
@@
Together with the jump in MSV, this bistability is suggestive of a discontinuous transition to active turbulence. To probe this scenario, we test whether the transition displays hysteresis. To this end, we take a single realization and progressively increase activity over time, letting the system relax to the new state after each activity increment (Section A.1). In the example in Fig. 1, where we observe a transition to turbulence around $A\approx 3900$ (red points). Following the reverse protocol, progressively decreasing activity from the turbulent state, the transition back to laminar flow happens at around $A\approx 3600$ (blue points in Fig. 1), showing a clear hysteresis cycle. The presence of hysteresis strongly supports the discontinuous nature of the transition to active turbulence.
III.2 Maximal Lyapunov exponent and the transition to chaos
@@
To show that this transition indeed corresponds to the emergence of chaos, we calculate the maximal Lyapunov exponent (MLE) $\Lambda$ (Section A.2). The MLE quantifies the global, long-time chaotic nature of a given initial condition by measuring the growth rate of a small deviation from it. For chaotic dynamics, the MLE is positive, while for regular, non-chaotic dynamics it is zero or negative 59. By computing the MLE of many realizations at different activities, we find a value of zero indicating non-chaotic dynamics at low activity, and a non-zero value indicating chaos at high activity (Fig. 1). In the chaotic state, the MLE increases linearly with activity. This trend is consistent with the experimental measurements by Tan et al. 24, which reported that, when multiplied by the time scale $\tau_{\text{v}}\equiv\ell_{\text{c}}/\sqrt{\langle v^{2}\rangle}\sim 1/A$, the MLE was independent of activity. Finally, in the transition region, we find both zero and positive MLE values. The presence of intermediate MLE values — neither zero nor corresponding to the chaotic regime — indicates that there are many realizations in the transition regime which have extremely long chaotic transients before finding a laminar state, and consequently the numerically estimated MLE has not yet decreased to zero.
@@
Given that the transition to active turbulence involves bistability of laminar and turbulent states, we consider the fraction of time that the system spends in the chaotic state $f_{c}$ as an order parameter (Fig. 1). This fraction can be constructed from the so-called stretching numbers or finite-time Lyapunov exponents $\alpha$ of each realization. Whereas the long-time MLE $\Lambda$ measures the average growth rate of a perturbation over the entire time of a realization, the stretching number $\alpha$ measures it over short time windows (Section A.1). Thus, positive values of $\alpha$ indicate chaotic periods while zero or negative values indicate non-chaotic periods.
@@
We obtain time series of $\alpha$, as exemplified in Fig. 1, which show clear switches from chaotic to laminar flow. We then build histograms of these time series (Fig. 1), which indicate the probability of finding the system in either the laminar or the chaotic state. At very low activity, the histogram has a peak at $\alpha=0$, indicating that the system is always in the laminar state (Fig. 1). In the transition region, realizations display either chaos for all their duration or a long chaotic transient that eventually decays to laminar flow (Fig. 1). At high activity, only chaos is observed (Fig. 1). Setting a long integration time ($6\times 10^{2}\ \tau_{\mathrm{r}}$), we separate the two peaks in the bimodal distribution and integrate their areas to obtain the chaotic fraction $f_{\text{c}}$, which transitions from zero at low activity to one at high activity (Fig. 1). Further work is required to understand the V-shaped behavior of the chaotic fraction in the transition region.
@@
At what value of the activity number does the transition to active turbulence occur? Chaotic transients become extremely long close to the transition (Fig. S1). As a result, identifying the transition point is limited by the finite integration time: States that are chaotic over a long simulation could still decay to laminar flow over longer times. To address this challenge, we perform simulations for increasingly long times. We find that, for $A\lesssim 4900$, the chaotic fraction $f_{\text{c}}$ goes down for longer simulations (Fig. 1, inset), which shows that chaotic states at these activities eventually laminarize. In contrast, at $A\gtrsim 4900$, the system remains chaotic at all times. Therefore, we identify the transition to turbulence at $A^{*}\approx 4900$.
@@
IV Sequence of bifurcations towards chaos
@@
Having studied the transition to active turbulence from a statistical physics perspective, we now turn to a dynamical systems approach. For a given realization, what sequence of bifurcations leads to chaos? To answer this question, we build the bifurcation diagram by simulating a single realization, starting from a uniformly aligned nematic at zero activity and progressively increasing the activity number over time (Fig. 2). As before, we use the MSV $\langle v^{2}\rangle_{\mathbf{r}}$ to quantify the flow intensity (purple in Fig. 2), but we now also consider its $x$ and $y$ components $\langle v_{x}^{2}\rangle_{\mathbf{r}}$ and $\langle v_{y}^{2}\rangle_{\mathbf{r}}$ (red and blue in Fig. 2) to distinguish between different flow patterns. To understand some of the rich dynamics, we also compute not only the largest, but the first six Lyapunov exponents (LEs) (Fig. 2, see methods in Section A.2).
IV.1 Onset of spontaneous flows and vortices through two supercritical pitchfork bifurcations
The first bifurcation takes place at $A\approx 100$, where a spontaneous simple shear flow (Fig. 2) emerges from the quiescent, no-flow state (dark to light grey regions in Fig. 2). This process corresponds to the well-known spontaneous-flow instability 60, which is a supercritical pitchfork bifurcation 61. Because we consider a contractile system with the director initially aligned along $\hat{\bm{x}}$, the spontaneous flow is also along $\hat{\bm{x}}$. Therefore, the flow breaks translational symmetry along $\hat{\bm{y}}$, which results in a Goldstone mode showing up as a new zero LE (orange points in Fig. 2). Note that, due to the breaking of rotational symmetry in the initial aligned state, there is one zero LE already at zero activity.
@@ IV.1 Onset of spontaneous flows and vortices through two supercritical pitchfork bifurcations
IV.2 First subcritical bifurcation and the onset of oscillations
Further increasing activity, at $A\approx 941$ we observe a first subcritical bifurcation 61, where both $\langle v_{x}^{2}\rangle_{\mathbf{r}}$ and $\langle v_{y}^{2}\rangle_{\mathbf{r}}$ jump discontinuously (green to red regions in Fig. 2). The system switches from a two-vortex to a three-vortex state (Figs. 2 and 2). These two states are bistable in a narrow range of activity, where we observe hysteresis (overlapping green-red region in Fig. 2).
@@ IV.2 First subcritical bifurcation and the onset of oscillations
Overall, these results are consistent with the oscillations reported in previous work 62; 45; 46; 48; 25. Here, our analysis shows that oscillations in active nematics emerge at the first subcritical bifurcation of vortical flows — far from the primary spontaneous-flow instability, which lacks oscillations 60; 55; 63.
IV.3 Second subcritical bifurcation and the onset of chaotic transients

### v2

@@ Abstract
Active fluids exhibit chaotic flows at low Reynolds number known as active turbulence. Whereas the statistical properties of the chaotic flows are increasingly well understood, the nature of the transition from laminar to turbulent flows as activity increases remains unclear. Here, through simulations of a minimal model of unbounded and defect-free active nematics, we find that the transition to active turbulence is discontinuous. We show that the transition features a jump in the mean-squared velocity, as well as bistability and hysteresis between laminar and chaotic flows. From distributions of finite-time Lyapunov exponents, we identify the transition at a value $A^{*}\approx 4900$ of the dimensionless activity number. Below the transition to chaos, we find subcritical bifurcations that feature bistability of different laminar patterns. These bifurcations give rise to oscillations and to chaotic transients, which become very long close to the transition to turbulence. Overall, our findings contrast with the continuous transition to turbulence in channel confinement, where turbulent puffs emerge within a laminar background. We propose that, without confinement, the long-range hydrodynamic interactions of Stokes flow suppress the spatial coexistence of different flow states, and thus render the transition discontinuous.
@@
How do laminar flows become turbulent? Scientists have been seeking answers to this deceptively simple question since the seminal experiments by Reynolds in 1883 1. Reynolds’ original work showed that the onset of turbulence is controlled by a dimensionless parameter — the Reynolds number — which compares inertial to viscous forces. Despite having a single control parameter, the nature of the transition to turbulence has been debated for over a century 2. Detailed experiments and simulations in the past decade showed that, in pipe and Couette flow, the onset of turbulence follows a continuous second-order phase transition, with critical exponents in the directed percolation universality class 3, 4, 5, 6, 7, 8, 9, 10, 11, 12. Even more recent work showed that, under body forces, the transition becomes discontinuous 13, 14, similar to first-order phase transitions.
@@
In stark contrast to high-Reynolds-number inertial turbulence, two decades ago, turbulent-like flows were discovered in bacterial suspensions at low Reynolds number 15. Such spontaneous chaotic flows, now broadly known as active turbulence, have since been observed in a variety of (predominantly biological) fluids 16. In addition to bacterial suspensions 15, 17, 18, 19, 20, 21, 22, 23, active turbulence is found in suspensions of microtubules and molecular motors 24, 25, 26, 27, 28, 29, 30, as well as in epithelial monolayers 31, 32, 33 and self-propelled particles 34. These active systems consume stored energy to power local internal driving 35. When this internal driving is strong enough, the flows become chaotic. But how and when does chaos emerge?
@@
For active polar fluids, simulations of multiple models showed that chaos is reached through a sequence of oscillatory instabilities 36, 37, 38, 39, 16, 40, consistently with recent experiments on bacterial suspensions 41. Other works have reported transitions to turbulence by varying either activity 42, 43, 44, 45 or swimmer concentration 46, 47, 23, 48. For active nematics, simulations showed that, in a channel, the system goes from simple shear to oscillatory flow and to a vortex chain before reaching turbulence49, 50, 51, 52, 53, 54. Similar states were seen experimentally in bacterial and in microtubule suspensions by increasing the confinement size 55, 27, 56. Importantly, simulations showed that the transition to chaos follows a scenario of directed percolation similar to pipe flow, with localised puffs of turbulence eventually spreading throughout the system as activity increases 51. In the absence of confinement, both simulations 57, 58, 59 and experiments 28 revealed a sequence of bend instabilities leading to turbulence. However, key questions remain: What is the nature of the bifurcations that end up in chaos? And what is the activity threshold, as well as the type, of the transition to turbulence?
@@
Here, we combine approaches from dynamical systems and statistical physics to investigate the transition to active nematic turbulence in a minimal model without confinement. We find that the transition is discontinuous, with the system exhibiting a jump in flow intensity and fluctuations, as well as a region of bistability and hysteresis, as in first-order phase transitions. We then build the bifurcation diagram of the initial instabilities that ultimately lead to the outbreak of chaos. Thereby, we expose the earliest appearances of subcritical bifurcations, oscillations, and chaotic transients that culminate in the discontinuous transition to turbulence. When compared to the directed-percolation scenario in channels 51, our findings suggest that confinement can change the nature of the transition from discontinuous to continuous. Therefore, our results limit the applicability of the notion of universality in the transition to active turbulence. Our results also imply that, when unconstrained, active nematic turbulence sets in at a value $A^{*}\approx 4900$ of the dimensionless activity number, which therefore plays a role similar to that of the critical Reynolds number for the transition to inertial turbulence.
Active nematics hydrodynamics
@@
We study a minimal hydrodynamic model of incompressible active nematics in two dimensions 59. Neglecting inertia by taking zero Reynolds number, force balance reads
$0=-\partial_{\alpha}P+\partial_{\beta}\left(\sigma_{\alpha\beta}+\sigma^{\text{a}}_{\alpha\beta}\right).$ (1)
@@
Here, the symmetric part of the stress tensor is $\sigma_{\alpha\beta}=2\eta v_{\alpha\beta}-\zeta q_{\alpha\beta}$, with $\eta$ the shear viscosity, $v_{\alpha\beta}=1/2\left(\partial_{\alpha}v_{\beta}+\partial_{\beta}v_{\alpha}\right)$ the symmetric part of the strain rate tensor, $\zeta$ the active stress coefficient, and $q_{\alpha\beta}=n_{\alpha}n_{\beta}-1/2\delta_{\alpha\beta}$ the nematic orientation tensor defined by the director field $\bm{n}=(\cos\theta,\sin\theta)$. The director has unit norm as we consider the system to be deep in the nematic phase and consequently defect-free. Respectively, the antisymmetric stress is $\sigma^{\text{a}}_{\alpha\beta}=1/2\left(n_{\alpha}h_{\beta}-h_{\alpha}n_{\beta}\right)$, where the molecular field $h_{\alpha}=K\nabla^{2}n_{\alpha}$ arises from minimizing the Frank free energy for nematic elasticity in the one-constant approximation 60.
@@
where $\omega_{\alpha\beta}=1/2\left(\partial_{\alpha}v_{\beta}-\partial_{\beta}v_{\alpha}\right)$ is the vorticity tensor and $\gamma$ is the rotational viscosity. For simplicity, following Ref. 59, we ignore flow alignment; including it does not alter the scaling properties of the turbulent flows 61.
@@
$\nabla^{4}\psi=\frac{1}{2}\frac{R}{A}\nabla^{4}\theta+S\left[\frac{1}{2}\left(\partial_{x}^{2}-\partial_{y}^{2}\right)\sin(2\theta)-\partial_{xy}^{2}\cos(2\theta)\right],$ (3)
where $S=\zeta/|\zeta|$ is the sign of the active stress, $R=\gamma/\eta$ is the viscosity ratio, and the activity number $A=L^{2}/\ell^{2}_{c}$ where $\ell_{c}=\sqrt{K/(|\zeta|R)}$ is the active length from balancing active and nematic elastic stresses. Due to the absence of flow alignment, contractile (S=1S=-1) and extensile ($S=1$) active stresses are equivalent in our model 62, 63, 64. In the following sections, we set $R=1$ and vary $A$. Since the active time $\tau_{\text{a}}$ itself varies with activity, we use the nematic relaxation time $\tau_{\text{r}}=\gamma L^{2}/K$ as the time unit in the plots.
@@
Equations 3 and 4 provide a minimal hydrodynamic model that captures the scalings of the velocity power spectrum of fully-developed turbulent flows 59, which have been successfully compared to experiments 30.
Transition to active turbulence
@@
Figure 1: Transition to active turbulence. 1-1, Snapshots of the flow field in the laminar (1), bistable (1,1), and chaotic (1) regimes. The background shows the nematic director through line integral convolution, color denotes the stream function $\psi$, and arrows depict the velocity. 1, Mean squared velocity (MSV) $\langle v^{2}\rangle_{\mathbf{r}}$ from multiple realizations at different activity number AA. In panels 1,$1,$ gray points are individual realizations, colored points are the ensemble averages, and the error bars are s.e.m. Regimes of laminar and chaotic flow are separated by a jump in MSV and a transition region where both states are possible. Labels 1-1 correspond to the snapshots above. 1, MSV from a single realization with increasing (red) or decreasing (blue) activity. The transition between laminar and turbulent flows exhibits hysteresis. Shaded regions show the s.d. of temporal fluctuations in MSV. 1, The maximal Lyapunov exponent (MLE) $\Lambda$ shows the transition from non-chaotic to chaotic dynamics as activity increases. 1, The fraction of time spent in the chaotic state $f_{\text{c}}$ also shows a transition from non-chaotic to chaotic dynamics with an intermediate region where both are possible. The inset zooms into the end of the transition region. By increasing the integration time $t_{\text{f}}$ (blue to red), we identify the transition to turbulence at $A^{*}\approx 4900$, above which the flow never laminarizes. 1, Exemplary time series of the stretching number $\alpha$ showing a chaotic transient ($\alpha\neq 0$) that eventually decays into a laminar state ($\alpha=0$). 1, Histogram of 1. 1-1, For increasing activity, the distributions of the stretching number show the change from completely laminar (1) to completely chaotic flow (1) through a transition region with a bimodal distribution indicative of bistability (1). The unit of time for all quantities in this figure is $\tau_{\text{r}}$.
@@
When and how do turbulent flows emerge? We first approach this question statistically. To find and characterize the transition to turbulence, we numerically integrate Eqs. 3 and 4 at different values of the activity number $A$, starting from a uniformly aligned director field (see Methods). To obtain different realizations, we add noise for a short initial period before the evolution continues deterministically (see Methods for details). Looking at many such realizations, we identify three broad regimes with qualitatively different dynamics: (i) At low activity, we find laminar flow states, such as vortex patterns, either steady or oscillating (Fig. 1 and Movie S1); (ii) at intermediate activity, both laminar flow patterns and chaos are possible (Figs. 1 and 1); and (iii) at high activity, we find chaotic flow (Fig. 1 and Movie S2).
Jump in flow intensity, bistability, and hysteresis
@@
To quantify the flow intensity in these different regimes, we compute the space-averaged mean squared velocity (MSV), $\langle v^{2}\rangle_{\mathbf{r}}$, for many realizations at each activity (Fig. 1). At low activity ($A\lesssim 3800$), the MSV increases linearly with activity. There is very little variation between realizations (grey points in Fig. 1; purple points show the ensemble average). At high activity, ($A\gtrsim 5000$), the MSV once again increases linearly, consistent with previous findings 65, though now with a steeper slope. In addition, the variation between realizations is much higher than in the laminar states at low activity. In between, we find a transition region ($3800\lesssim A\lesssim 5000$) showing bistability between laminar and chaotic flows: Both types of flow can occur at a given activity.
@@
Together with the jump in MSV, this bistability is suggestive of a discontinuous transition to active turbulence. To probe this scenario, we test whether the transition displays hysteresis. To this end, we take a single realization and progressively increase activity over time, letting the system relax to the new state after each activity increment (Methods). In the example in Fig. 1, we observe a transition to turbulence around $A\approx 3900$ (red points). Following the reverse protocol, progressively decreasing activity from the turbulent state, the transition back to laminar flow happens at around $A\approx 3600$ (blue points in Fig. 1), showing a clear hysteresis cycle. The presence of hysteresis strongly supports the discontinuous nature of the transition to active turbulence.
Maximal Lyapunov exponent and the transition to chaos
@@
To show that this transition indeed corresponds to the emergence of chaos, we numerically estimate the maximal Lyapunov exponent (MLE) $\Lambda$ (Methods). The finite-time estimate of the MLE quantifies the global, long-time chaotic nature of a given initial condition by measuring the growth rate of a small deviation from it. For chaotic dynamics, the MLE is positive, while for regular, non-chaotic dynamics it is zero or negative 66. By computing the MLE of many realizations at different activities, we find a value of zero indicating non-chaotic dynamics at low activity, and a non-zero value indicating chaos at high activity (Fig. 1). In the chaotic state, the MLE increases linearly with activity. This trend is consistent with the experimental measurements by Tan et al. 26, which reported that, when multiplied by the time scale $\tau_{\text{v}}\equiv\ell_{\text{c}}/\sqrt{\langle v^{2}\rangle}\sim 1/A$, the MLE was independent of activity. Finally, in the transition region, we find both zero and positive MLE values. The presence of intermediate MLE values — neither zero nor corresponding to the chaotic regime — indicates that there are many realizations in the transition regime which have extremely long chaotic transients before finding a laminar state, and consequently the numerically estimated finite-time MLE has not yet decreased to zero.
@@
Given that the transition to active turbulence involves bistability of laminar and turbulent states, we consider the fraction of time that the system spends in the chaotic state $f_{c}$ as an order parameter (Fig. 1). This fraction can be constructed from the so-called stretching numbers or finite-time Lyapunov exponents $\alpha$ of each realization. Whereas the long-time MLE $\Lambda$ measures the average growth rate of a perturbation over the entire time of a realization, the stretching number $\alpha$ measures it over short time windows (Methods). Thus, positive values of $\alpha$ indicate chaotic periods while zero or negative values indicate non-chaotic periods.
@@
We obtain time series of $\alpha$, as exemplified in Fig. 1, which show clear switches from chaotic to laminar flow. We then build histograms of these time series (Fig. 1), which indicate the probability of finding the system in either the laminar or the chaotic state. At very low activity, the histogram has a peak at $\alpha=0$, indicating that the system is always in the laminar state (Fig. 1). In the transition region, realizations display either chaos for all their duration or a long chaotic transient that eventually decays to laminar flow (Fig. 1). At high activity, only chaos is observed (Fig. 1). Setting a long integration time ($6\times 10^{2}\ \tau_{\mathrm{r}}$), we separate the two peaks in the bimodal distribution and integrate their areas to obtain the chaotic fraction $f_{\text{c}}$, which transitions from zero at low activity to one at high activity (Fig. 1).
@@
At what value of the activity number does the transition to active turbulence occur? We define the transition to take place at an activity $A^{*}$ beyond which only chaos is observed, so that the chaotic fraction is one: $f_{\text{c}}(A>A^{*})=1$. This criterion identifies when chaos becomes persistent. However, chaotic transients become extremely long close to the transition (Fig. S1). As a result, identifying the transition point is limited by the finite integration time: States that are chaotic over a long simulation could still decay to laminar flow over longer times. To address this challenge, we perform simulations for increasingly long times. We find that, for $A\lesssim 4900$, the chaotic fraction $f_{\text{c}}$ goes down for longer simulations (Fig. 1, inset), which shows that chaotic states at these activities eventually laminarize. In contrast, at $A\gtrsim 4900$, the system remains chaotic at all times. Therefore, we identify the transition to turbulence at $A^{*}\approx 4900$.
Finally, right past the transition to turbulence, we find that at least six Lyapunov exponents are positive (Fig. S2). Thus, the transition seems to be directly to high-dimensional chaos.
@@
Sequence of bifurcations towards chaos
@@
Having studied the transition to active turbulence from a statistical physics perspective, we now turn to a dynamical systems approach. For a given realization, what sequence of bifurcations leads to chaos? To answer this question, we build the bifurcation diagram by simulating a single realization, starting from a uniformly aligned nematic at zero activity and progressively increasing the activity number over time (Fig. 2). As before, we use the MSV $\langle v^{2}\rangle_{\mathbf{r}}$ to quantify the flow intensity (purple in Fig. 2), but we now also consider its $x$ and $y$ components $\langle v_{x}^{2}\rangle_{\mathbf{r}}$ and $\langle v_{y}^{2}\rangle_{\mathbf{r}}$ (red and blue in Fig. 2) to distinguish between different flow patterns. To understand some of the rich dynamics, we also compute not only the largest, but the first six Lyapunov exponents (LEs) (Fig. 2, see methods in Methods).
Onset of spontaneous flows and vortices through two supercritical pitchfork bifurcations
The first bifurcation takes place at $A\approx 100$, where a spontaneous simple shear flow (Fig. 2) emerges from the quiescent, no-flow state (dark to light grey regions in Fig. 2). This process corresponds to the well-known spontaneous-flow instability 67, which is a supercritical pitchfork bifurcation 68. Because we consider a contractile system with the director initially aligned along $\hat{\bm{x}}$, the spontaneous flow is also along $\hat{\bm{x}}$. Therefore, the flow breaks translational symmetry along $\hat{\bm{y}}$, which results in a Goldstone mode showing up as a new zero LE (orange points in Fig. 2). Note that, due to the breaking of rotational symmetry in the initial aligned state, there is one zero LE already at zero activity.
@@
First subcritical bifurcation and the onset of oscillations
Further increasing activity, at $A\approx 941$ we observe a first subcritical bifurcation 68, where both $\langle v_{x}^{2}\rangle_{\mathbf{r}}$ and $\langle v_{y}^{2}\rangle_{\mathbf{r}}$ jump discontinuously (green to red regions in Fig. 2). The system switches from a two-vortex to a three-vortex state (Figs. 2 and 2). These two states are bistable in a narrow range of activity, where we observe hysteresis (overlapping green-red region in Fig. 2).
@@
Overall, these results are consistent with the oscillations reported in previous work 69, 49, 50, 52, 27. Here, our analysis shows that oscillations in active nematics emerge at the first subcritical bifurcation of vortical flows — far from the primary spontaneous-flow instability, which lacks oscillations 67, 59, 64. Our finding that oscillations emerge from a vortex state in unconfined active nematics contrasts with the results of previous simulations in channel confinement, where oscillatory flow appeared before vortices 52. These variations highlight the influence of boundary conditions on the bifurcation diagram of active nematics.
Second subcritical bifurcation and the onset of chaotic transients